OpenStudy (anonymous):

Find the slope-intercept equation of the line that has the given characteristics. Slope 0 and y-intercept (0,9)

OpenStudy (anonymous):

well you have a point (0,9) and a slope (0). Write the equation in point-slope form and then solve for y.

OpenStudy (anonymous):

So it would be 9=0(0)+b

OpenStudy (anonymous):

point-slope is y-y1=m(x-x1), so the point-slope form would be y-9=0(x-0), then you would solve for y.

OpenStudy (anonymous):

Y would be 9.

OpenStudy (anonymous):

yep, the equation is y=9 :)